please walk throughLets modify logsic differantal equation of this example as followa 0.05P1- - ..natnn nf fish s . Tiithe r i Eh are caught (b this for me.
Answer #1

a) (-8) represents harvesting the fish at weekly rate. 8 fish/week is the rate at which fish are caught

b) Direction field is plotted below

1200 1000 800 600 400 200 0 0 10 20 40 50 60 70 80 90 30 100

Which corresponds to the last option

c) Equilibrium solutions are: P(t)200,800 (based on the above graph where the derivative is horizontal)

d) Graph of derivative function is:

10 f(x0.05.x| 1-1000 (800, 0) (200, 0) 200 800 0 400 600 1000 1200 10

And so

0 Po<200 P(t) is decreasing Po 200 P(t) is fixed 200 < Po800 P(t) is increasing Po 800 P(t) is fixed Po> 800 P(t) is decreasi

d) General solution to a) is y(t)=200-\frac{600}{13\cdot\exp\left(-\frac{3t}{100}\right)-1}

General solution to b) is y(t)=200+\frac{600}{11\cdot\exp\left(-\frac{3t}{100}\right)+1}

The two solutions are plotted below:

+ -500 X f(x)= 0.05x1l - 1000 -8 400 600 N g(x) 200- = -1 13 exp 100 300 X 600 Nh(x) =200+ 3x 11 exp 100 200 4 100 400 300 20

Which corresponds to option c) (the lower left option)

\blacksquare

Please do rate this answer positively if you found it helpful. Thanks and have a good day!